OpenStudy (anonymous):

Find the vertex and the axis of symmetry of the graph of the function: f(x) = -2x2 – 4x?

OpenStudy (anonymous):

The first step would be to find the vertex. Do you know how to write a given parabola in vertex form?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

One way is to complete the square. It allows you to rewrite the standard form \(ax^2+bx+c\) into vertex form \((x-h)^2+k\), where \((h,k)\) is the vertex. So we're given \[y=-2x^2-4x\] Do you know how to complete the square?

OpenStudy (anonymous):

is it -1 and-4?

OpenStudy (anonymous):

aaa or -6 -4

OpenStudy (anonymous):

\[\begin{align*}y&=-2x^2-4x\\ &=-2\left(x^2+2x\right)\\ &=-2\left(x^2+2x+1-1\right)\\ &=-2\left((x+1)^2-1\right)\\ &=-2(x+1)^2+2\\ \end{align*}\] What's the vertex?

OpenStudy (anonymous):

lost because is not the answer in the back of my book

OpenStudy (anonymous):

Well, according to my work, the vertex is \((-1,2)\), and Wolfram agrees: http://www.wolframalpha.com/input/?i=vertex+of+-2x%5E2-4x

OpenStudy (anonymous):

sorry looks like i didnt put this at the end of the question What are the x-intercepts of y = f(x-7

OpenStudy (anonymous):

of \(y=f(x-7)\) or \(f(x)-7\) ?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Okay. If \(f(x)=-2x^2-4x\), then what is \(f(x-7)\) ?

OpenStudy (anonymous):

no clue lol

OpenStudy (anonymous):

Plug in \(x-7\): \[\begin{align*}f(x)=-2x^4-4x~~\Rightarrow~~f(x-7)&=-2(x-7)^2-4(x-7)\\ &=-2\left(x^2-14x+49\right)-4x+28\\ &=-2x^2+28x-98-4x+28\\ &=-2x^2+24x-70\end{align*}\] Make sense so far?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

the problem is thisFind the vertex and the axis of symmetry of the graph of the function: f(x) = -2x^2 – 4x

OpenStudy (anonymous):

and this are the answers (-1,0); x = -1 (-1,2); x = -1 (1,2); x = 1 (1,0); x = 1

OpenStudy (anonymous):

which am still not sure off

OpenStudy (anonymous):

Hmm, something's not working with the equation editor. Give me a moment to write it out.

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

 So what's the vertex?

OpenStudy (anonymous):

1-1

OpenStudy (anonymous):

Hmm? You should be getting a vertex of \((6,2)\)... http://www.wolframalpha.com/input/?i=vertex+of+-2%28x-7%29%5E2-4%28x-7%29

OpenStudy (anonymous):

mmm

OpenStudy (anonymous):

i think the book is confusing me

OpenStudy (anonymous):

Answer key could be mismatched with a different question.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks though

OpenStudy (anonymous):

yw